• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 11:06:52 on February23,2001

    Entry number 56837

    keyword=He4 re-cool-down completed

    He4 target re-cooldown (with cold-return) started by Joe, was completed
    at 10:45 am. The condition now is at 5.8K and 212 psia (read by the tranducer,
    the gauge at the gas panel reads 222 psia). Fan is at 25% and HP heater is in PID
    with about 50 watts and 4K flow is 34 g/s. Coolant temperatures: ESR supply
    is 4.25K, T after JT is 4.6, after HX is 6.4K, and on the target return line is
    7.85K and after mixed with spectromter return, it is 5.25K (this one normally should
    be below 5K). So it is still a little high. Will take some time to settle.